Lou has quarters and dimes in his own coin bank. There are 26 coins in the bank. If he add 7 quarters and 3 dimes, the total value of the coins will be $7.35. Find the no. of quarters and dimes Lou originally have.
Answer by AnlytcPhil(1806)

Lou has quarters and dimes in his own coin bank. There
are 26 coins in the bank. If he add 7 quarters and 3 
dimes, the total value of the coins will be $7.35. 
Find the no. of quarters and dimes Lou originally have.

Q + D = 26
.25(Q + 7) + .10(D + 3) = 7.35 

Simplify the second equation and solve the system.

Answer: 18 quarters and 8 dimes.

Check:  18+8 = 26 coins

18+7 or 25 quarters is worth $6.25
8+3 or 11 dimes is worth $1.10
That would be a total of $7.35

Edwin